OpenStudy (anonymous):

MEDALS COME AND GET THEM !!!!!!!!!!!!! A line has a slope of 7 and contains the point (5, −9). Which equations represent the line? Choose all answers that are correct. A. −7x + y = −44 B. 7x – y = 44 C. (y − 9) = 7(x + 5) D. (y + 9) = 7(x − 5)

OpenStudy (anonymous):

who ever can help me answer the most gets the medals

OpenStudy (anonymous):

@iambatman

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

Do you have an idea of the answer?

OpenStudy (anonymous):

@anyone

OpenStudy (anonymous):

no i am completely stuck and i feel like a dumb butt ;-;

Nnesha (nnesha):

y=mx+b this is slope intercept form equation where m is slope which is always with x and b is y-intercept so replace m by 7

Nnesha (nnesha):

then they gave you one order pair where first values is x and 2nd is y like this (x,y) <-- plug that values into x and y and then solve for b

OpenStudy (anonymous):

can you give an example its hard for me to grasp to the process

OpenStudy (anonymous):

5=x -9=y Try to fill on the the x and y for each answer and see what works.

Nnesha (nnesha):

\[\huge\ y-y _{1}=m(x-x _{1})\] point slope form

OpenStudy (anonymous):

ok ill try it dont leave yet

Nnesha (nnesha):

use that one ^^^

Nnesha (nnesha):

hha don't worry take ur time ii'll be glad to wait :)

OpenStudy (anonymous):

is it C and D?

OpenStudy (anonymous):

im sorry its really late for me and im not rolly tinking straight

Nnesha (nnesha):

 replace variables by their values